OpenStudy (anonymous):

line AB is tangent to circle C at point B...Which of the following is not true?

OpenStudy (anonymous):

\[\angle CAB < \angle ACB\] \[AB^2 + BC^2 = AC^2 \] \[\angle CAB and \angle ACB are complements\] \[Line AB is perpendicular \to line BC\]

OpenStudy (anonymous):

These are the options on my review

OpenStudy (epoweritheta):



OpenStudy (anonymous):

so B is true...

OpenStudy (anonymous):

and A

OpenStudy (epoweritheta):

I assume that the centre of the circle is C . (a) is not necessarily true . (b) is always true (rt angle) (c)true (since they sum up to 90) (d)true (since tangent at any point is perpendicular to radius through that point )

OpenStudy (epoweritheta):



OpenStudy (anonymous):

oh...so for A it could be a smaller angle sometimes?

OpenStudy (epoweritheta):

see the above picture they cannot be compared

OpenStudy (anonymous):

oh ok that makes sense....I get why D is true...but how do u know they are complements for C?

OpenStudy (epoweritheta):

A is complement to B if A+B=90 , here both those angles add up to 90 , because the 3rd angle is 90 and sum of angles in a triangle is 180

OpenStudy (anonymous):

Got it! thank you!!!